A certain volume of gas is found to weigh 4.4g, an equal volume of hydrogen gas weighs 0.2g. The molar mass of the gas is


1
Expert's answer
2021-06-11T05:37:38-0400

The volume of the gases is equal, thus number of moles as well:

"V_x = V_{H_2} \\
n_x \times V_m = n_{H_2} \times V_{H_2}\\
n_x = n_{H_2}\\"


Let's find the number of moles for the hydrogen gas

"n_{H_2} = \frac{m_{H_2}}{M_{H_2}} = \frac{0.2 g}{2 g/mol}=0.1 mol\\"


Then, using the values of the n for the hydrogen, the molar mass of a gas can be derived:

"M_x=\frac{m_x}{n_{H_2}} = \frac{4.4 g}{0.1 mol}= 44 g/mol"